Circular No. 9040

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

2002 VF_{130}

K. S. Noll and S. D. Benecchi, Space Telescope Science Institute; and W. M. Grundy, Lowell Observatory, report the detection of a binary companion to the transneptunian object 2002 VF_{130} (MPEC 2002-W30, 2004-V71). The observations were made during 2008 Nov. 3.8076–3.8396 UT with the Wide Field Planetary Camera 2 on the Hubble Space Telescope, using the F606W filter (wide V) for four dithered 260-s exposures. The two components were widely separated, with an angular distance of 0′′.733 ± 0′′.003. The secondary binary component was fainter by 0.31 magnitude. The secondary was located at −0′′.576 ± 0′′.003 in α and 0′′.453 ± 0′′.001 in δ relative to the primary.

COMET P/2009 H2 (McNAUGHT) Automatic procedures at the Minor Planet Center identified comet

P/2004 K2 (cf. IAUC 8348) in incidental astrometry obtained by T. H. Bressi with the Spacewatch 0.9-m f/3 reflector on Apr. 28. Subsequently, recovery observations were independently reported by G. Muler, J. M. Ruiz, and R. Naves with the 0.30-m Schmidt-Cassegrain telescope at the Observatorio Nazaret (Lanzarate, Spain) on May 1 and 3. No information was provided about the comet’s appearance. The first observation of each set is provided below:

2009 UT α_{2000} δ_{2000} Mag. Observer
Apr. 28.36557 16^{h}03^{m}06.^{s}17 −7^{o}33′49′′.0  Spacewatch
May 1.04433 16 01 23.07 −7 18 04.3 20.2 Muler
3.05734 15 59 59.04 −7 06 17.5 20.0  _{”}

The indicated correction to the predictions on MPC 56805 and in the 2008–2009 Comet Handbook is ΔT = −0.08 day. The remaining astrometry and the following (and additional) orbital elements from 155 observations spanning 2004 May 20–2009 May 3, together with residuals and an ephemeris, appear on CBET 1790.

Epoch = 2010 Jan. 4.0 TT

T = 2009 Dec. 15.4321 TT  ω = 180.^{o}7575
e = 0.502681  Ω = 150.1181 }2000.0
q = 1.548625 AU  i = 8.1327

a = 3.113947 AU n^{o} = 0.1793648 P = 5.495 years

2009 May 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green